OpenStudy (anonymous):

how do u find the domain?? g(x)= sqrx+5 ??

OpenStudy (anonymous):



OpenStudy (accessdenied):

The domain is the set of numbers \(x \) that you can put into \(g(x)\) to get a real output. Consider the parts of your function that may have values that make the expression undefined or not real and set up an inequality for it (or not-equal-to).

OpenStudy (anonymous):

so that would just be -5?

OpenStudy (accessdenied):

\[ \large g(x) = \underbrace{\sqrt{x+5}}_{nonnegative}\\ \implies x + 5 \ge 0 \]

OpenStudy (accessdenied):

Well, you'll have -5 as one of the points in the domain (-5+5 = 0, sqrt(0) = 0 exists), but any number greater than -5 will also be in the domain because they also work.

OpenStudy (anonymous):

then -5 to infinity?

OpenStudy (accessdenied):

So, the domain would be all real numbers \(x \ge -5\), or \([-5,\infty)\), yeah.